# Grand-Canonical Typicality

### v1

@@ 1 Introduction
in the micro-canonical and canonical case, where $u_{\mathbb{S}}$ means the normalized uniform distribution over the sphere $\mathbb{S}$ (i.e., the surface area measure such that $u_{\mathbb{S}}(\mathbb{S})=1$) and $\mathrm{GAP}_{\hat{\rho}}$ the GAP measure [22, 16, 14, 44, 27, 50, 28] with density matrix $\hat{\rho}$ (i.e., the most spread-out measure with density matrix $\hat{\rho}$, see Section 7.1 for the definition). Among other things, we determine $\mathbb{P}_{\mathrm{gc}}$ in this paper (see Statements 3 and 4b in Section 3).
@@
under the constraint that the expected energy is fixed, $\operatorname{tr}(\hat{\rho}\,\hat{H})=E$ (as well as, in the grand-canonical case, the expected particle numbers, $\operatorname{tr}(\hat{\rho}\,\hat{N}_{i})=n_{i}$). While this reasoning may suggest that $\hat{\rho}_{\mathrm{can}}$ ($\hat{\rho}_{\mathrm{gc}}$) represents the knowledge of an observer who knows not more about the state than $E$ (and $n_{i}$) (e.g., [21, 54], [1, Chap. 4], [2, Chap. 7]) the reasoning outlined above in terms of $S\cup B$ provides more [31]: a physical reason why the actual, physical state (of $S$) is given (approximately) by $\hat{\rho}_{\mathrm{can}}$ (or $\hat{\rho}_{\mathrm{gc}}$); see Section 4.1 for further discussion.
Also concerning the use of the micro-canonical distribution $u_{\mathrm{mc}}=u_{\mathbb{S}(\mathscr{H}_{\mathrm{mc}})}$ or density matrix $\hat{\rho}_{\mathrm{mc}}$, there are different attitudes (e.g., [2, Sec. 6.8]) in the literature: Sometimes its use is regarded as a “principle of equal a priori probabilities” (e.g., [45, 20, 30]) saying that unless we have further knowledge about $\Psi$ than that it lies in $\mathscr{H}_{\mathrm{mc}}$, it is a rational guess to assign each point on $\mathbb{S}(\mathscr{H}_{\mathrm{mc}})$ equal (subjective) probability. Our attitude is different: Since it is the nature of thermal equilibrium to behave like most $\Psi$, we find the thermal equilibrium behavior by studying the behavior of most $\Psi$, where “most” refers to $u_{\mathrm{mc}}$ or, when appropriate, $u_{\mathrm{gmc}}$.
@@
$\hat{\rho}_{\mathrm{gG}}=Z^{-1}_{\mathrm{gG}}\;\exp\biggl(\sum_{k=1}^{K}\lambda_{k}\hat{Q}_{k}\biggr)\,$, (7)
which we also discuss as it plays, in fact, a key role. It represents the thermal equilibrium ensemble for which $\hat{Q}_{1},\ldots,\hat{Q}_{K}$ (and only those) are conserved macroscopic observables, [^1] The expression “macroscopic observable” is often intended to include the assumption that the eigenvalues of $\hat{Q}_{k}$ are separated by the resolution of macroscopic measurements; this is not intended here. assuming they commute with each other. (One of them usually is the Hamiltonian $\hat{H}$. Previous considerations in this direction can be found in [25, Sec. I.4].) Note that, as a consequence, there is a $K$-parameter family of thermal equilibrium states with parameters $\lambda_{1},\ldots,\lambda_{K}$ (which includes the canonical case with $K=1$, $\hat{Q}_{1}=\hat{H}$, and $\lambda_{1}=-\beta$). The generalized micro-canonical subspace $\mathscr{H}_{\mathrm{gmc}}$ for this situation is the one where the eigenvalues of each $\hat{Q}_{k}$ are restricted to an interval $Q_{k}-\Delta Q_{k},Q_{k}]$ that is short on the macroscopic scale but still has a high-dimensional spectral subspace $\mathds{1}_{[Q_{k}-\Delta Q_{k},Q_{k}]}(\hat{Q}_{k})\mathscr{H}$,
@@
We can give here a derivation in the case in which each of the $\hat{Q}_{k}$ is (approximately) extensive, that is, in which, for spatial regions $S$,
$\hat{Q}_{k}\approx\hat{Q}_{k}^{S}\otimes\hat{I}^{S^{c}}+\hat{I}^{S}\otimes\hat{Q}_{k}^{S^{c}}\,.$ (13)
This is the first of our claims, see Statement 1a in Section 3. It includes the grand-canonical density matrix $\hat{\rho}_{\mathrm{gc}}$ as a special case (also for chemical equilibrium, see Section 2).
@@ 2 Use of $\hat{\rho}_{\mathrm{gc}}$: Chemical Reactions
We answer both in this section.
@@
we assume that the interaction term $\hat{V}$ is mostly negligibly small (namely, except during a chemical reaction) and contains non-zero transition elements corresponding to every reaction (14), but not for any transition that is not of the form (14). That is, we assume that the conserved quantities of (14) are the only macroscopic conserved observables. In detail, using (simple) tools from linear algebra, consider the $r$-dimensional space $\mathcal{R}$ with axes labeled $n_{1},\ldots,n_{r}$, and let $\mathcal{L}$ be the subspace spanned by the $L$ vectors $(\nu_{\ell 1}-\tilde{\nu}_{\ell 1},\ldots,\nu_{\ell r}-\tilde{\nu}_{\ell r})$. Any change in particle numbers due to reactions of the form (14) lies in $\mathcal{L}$. Let
$K=r+1-\dim\mathcal{L}$ (24)
(which is equal to r+1Lr+1-L if the $L$ vectors just mentioned are linearly independent), and choose a linear mapping (a $(K-1)\times r$-matrix) $F:\mathcal{R}\to\mathbb{R}^{K-1}$ with kernel
@@
$F$ has full rank, $\mathrm{rank}(F)=\dim\mathrm{image}(F)=K-1$, and for every $k=1,\ldots,K-1$, $F_{k}(n_{1},\ldots,n_{r})$ is conserved during every reaction (14). As a consequence, the observables
@@
using (19); the function $(1/\beta)\log Z_{\mathrm{gc}}(\beta,\mu_{01},\ldots,\mu_{0r})$ is known as the grand potential (e.g., [1, (5.73)], [39, (2.7.13)]). Note that, while (37c) is an expression that one might have guessed right from the start, our derivation has achieved two things: First, provided reasons for why (37) is true, and second, provided a way of computing the $\mu_{0i}$ from the known quantities $n_{0i},\delta E_{\ell}$ by means of (35) and (10).
@@
In the example of (6), $E_{0A}$ and $E_{0B}$ can be chosen arbitrarily, $E_{0C}$ is determined by (20), (35) reduces to $\mu_{0C}=\mu_{0A}+\mu_{0B}+\delta E$, and (10) reduces to the three conditions
@@ 3 Main Claims
We now present our main statements. Apart from Statement 1b, the derivations of the statements are given below in the following sections. Recall that $S^{c}=\Lambda\setminus S$ denotes the spatial complement of $S$. We take $\hat{\rho}_{\mathrm{gmc}}$ as defined in (9) and (8), and $\hat{\rho}_{\mathrm{gG}}$ as defined in (7). For extensive $\hat{Q}_{k}$ as in (13), it follows that
$\hat{\rho}_{\mathrm{gG}}\approx\hat{\rho}_{\mathrm{gG}}^{S}\otimes\hat{\rho}_{\mathrm{gG}}^{S^{c}}$ (39)
with $\hat{\rho}_{\mathrm{gG}}^{S}$ formed according to (7) from the $\hat{Q}_{k}^{S}$ and likewise for $S^{c}$. This also entails that
$\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gG}}\approx\hat{\rho}_{\mathrm{gG}}^{S}\,.$ (40)
  • Statement 1a. (Generalized Gibbs Density Matrix) Let $\mathscr{H}$ be a Hilbert space and $\Lambda$ a subset of $\mathbb{R}^{3}$ (called “the available volume”). Suppose for every $S\subseteq\Lambda$, $\mathscr{H}$ factorizes into $\mathscr{H}=\mathscr{H}^{S}\otimes\mathscr{H}^{S^{c}}$. Suppose the self-adjoint operators $\hat{Q}_{1},\ldots,\hat{Q}_{K}$ on $\mathscr{H}$ are bounded from below and commute with each other, and that each is (approximately) extensive as in (13). Then for $\lambda_{1},\ldots,\lambda_{K}$ satisfying (10) and small $S\subset\Lambda$,
    $\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}\approx\hat{\rho}_{\mathrm{gG}}^{S}\,.$ (41)
@@
    $\mathscr{H}_{1i}=\mathscr{H}_{1i}^{S}\oplus\mathscr{H}_{1i}^{S^{c}}$ (43)
@@
    $\hat{H}_{1i}\approx\hat{H}_{1i}^{S}\oplus\hat{H}_{1i}^{S^{c}}\,.$ (44)
    Suppose (14) for $\ell=1,\ldots,L$ are the (only) possible reactions (i.e., changes in particle numbers), and (21) holds. Choose any $F$ satisfying (25), and define $\hat{Q}_{k}$ by (27) (so that $\hat{H},\hat{Q}_{k}]=0$) and $\hat{Q}_{K}=\hat{H}$. Then for $\mu_{0i}$ satisfying (10) and (35),
@@
    $\hat{\rho}_{\Psi}^{S}:=\operatorname{tr}^{S^{c}}|\Psi\rangle\langle\Psi|\approx\hat{\rho}_{\mathrm{gG}}^{S}\,.$ (46)
@@
    $\hat{\rho}_{\Psi}^{S}=\operatorname{tr}^{S^{c}}|\Psi\rangle\langle\Psi|\approx\hat{\rho}_{\mathrm{gc}}^{S}\,.$ (47)
@@
The conditional wave function can be thought of as what the collapsed wave function of $S$ would be after an ideal quantum measurement of the basis $b$ on $B$ [14, Footnote 2].
@@
We now answer the question, raised in the introduction, of what $\mathbb{P}_{\mathrm{gc}}$ is, the probability distribution of the wave function corresponding to the grand-canonical ensemble. We answer it by providing the typical distribution of the conditional wave function $\psi^{S}$. The answer is twofold, as it depends on how the basis $b$ is chosen. In general, $\psi^{S}$ (and its distribution) depends on the choice of $b$, and while $b$ can be chosen arbitrarily, one is often particularly interested in taking $b$ the position (or configuration) basis [4]. We provide two results assuming that $S$ is a small spatial region and $B=S^{c}$: one result in which $b$ is a typical (i.e., random) ONB, and one in which $b$ is typical among those that diagonalize the conserved operators $\hat{Q}_{k}^{S^{c}}$ (respectively, the particle number operators $\hat{N}_{i}^{S^{c}}$). The first result (Statement 3) applies whenever $b$ is unrelated to the joint eigenbasis of $\hat{H}^{S^{c}}$ and the $\hat{Q}_{k}^{S^{c}}$ (respectively, $\hat{N}_{i}^{S^{c}}$); the second (Statements 4a and 4b) whenever there is no simple relation between $b$ and $\hat{H}^{S^{c}}$ other than that both $b$ and a suitable eigenbasis of $\hat{H}^{S^{c}}$ diagonalize the $\hat{Q}_{k}^{S^{c}}$ (the $\hat{N}_{i}^{S^{c}}$)—this seems like a reasonable approximation for the position basis.
@@
  • Statement 3. ($\mathbb{P}_{\mathrm{gG}}$ and $\mathbb{P}_{\mathrm{gc}}$ for Typical ONB) In the setting of Statement 1a (or 1b) with $S$ a small spatial region, for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$ and most ONBs $b$ of $\mathscr{H}^{S^{c}}$, the conditional wave function (48) has distribution
@@
  • Statement 4a. ($\mathbb{P}_{\mathrm{gG}}$ for ONB Diagonalizing $\hat{Q}_{k}^{S^{c}}$) In the setting of Statement 1a, suppose that $S\subset\Lambda$ is a small region and that for every $k=1,\ldots,K-1$, $\Delta Q_{k}=0$ and $Q_{k}$ is an eigenvalue of $\hat{Q}_{k}$. Then for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$ and most ONBs $b$ of $\mathscr{H}^{S^{c}}$ that diagonalize $\hat{Q}_{1}^{S^{c}},\ldots,\hat{Q}_{K-1}^{S^{c}}$ (but not $\hat{Q}^{S^{c}}_{K}=\hat{H}^{S^{c}}$),
@@
  • Statement 4b. ($\mathbb{P}_{\mathrm{gc}}$ for ONB Diagonalizing $\hat{N}_{i}^{S^{c}}$) In the setting of Statement 1b, suppose that $S\subset\Lambda$ is a small region and that for every $k=1,\ldots,K-1$, $\Delta Q_{k}=0$ and $Q_{k}$ is an eigenvalue of $\hat{Q}_{k}$. Then for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$ and most ONBs $b$ of $\mathscr{H}^{S^{c}}$ that diagonalize $\hat{N}_{1}^{S^{c}},\ldots,\hat{N}_{r}^{S^{c}}$,
    $\mathscr{L}_{\psi^{S}}\approx\mathbb{P}_{\mathrm{gc}}:=\bigoplus_{q_{1}^{S},...,q_{K-1}^{S}}p(q_{1}^{S},\ldots,q_{K-1}^{S})~\mathrm{GAP}_{\hat{\rho}^{S}(q_{1}^{S},\ldots,q_{K-1}^{S})}\,$, (58)
@@
    $p(q_{1}^{S},\ldots,q_{K-1}^{S})=\operatorname{tr}(\hat{\rho}^{S}_{\mathrm{gc}}\hat{P})~~\text{with}~~\hat{P}:=\hat{P}\Bigl(\hat{Q}_{1}^{S}=q_{1}^{S},\ldots,\hat{Q}_{K-1}^{S}=q_{K-1}^{S}\Bigr)$ (59)
@@
Statement 4b is actually not a special case of Statement 4a because only certain combinations of $\hat{N}_{i}$ are conserved as the $\hat{Q}_{k}$, and while in Statement 4a we considered bases diagonalizing the $\hat{Q}_{k}^{S^{c}}$ (i.e., certain combinations of the $\hat{N}_{i}^{S^{c}}$), we consider in Statement 4b bases diagonalizing all $\hat{N}_{i}^{S^{c}}$. That is because we are interested in the position/configuration basis, which diagonalizes all $\hat{N}_{i}^{S^{c}}$; on the other hand, Statement 4a is the simpler statement, and can be applicable to general $\hat{Q}_{k}$ that are not related to number operators.
@@
The next statement concerns the approach (time evolution) towards thermal equilibrium. As in many earlier works such as [51, 32, 26, 12], we take “approach” to mean that for most $t\geq 0$ in the long run, the system is in thermal equilibrium. And we take “thermal equilibrium” to mean that $\hat{\rho}_{\Psi_{t}}^{S}\approx\hat{\rho}_{\mathrm{gG}}^{S}$ and $\mathscr{L}_{\psi^{S}}\approx\mathbb{P}_{\mathrm{gG}}$.

### v2

@@ 1 Introduction
in the micro-canonical and canonical case, where $u_{\mathbb{S}}$ means the normalized uniform distribution over the sphere $\mathbb{S}$ (i.e., the surface area measure such that $u_{\mathbb{S}}(\mathbb{S})=1$) and $\mathrm{GAP}_{\hat{\rho}}$ the GAP measure [25, 19, 17, 50, 31, 56, 32, 33] with density matrix $\hat{\rho}$ (i.e., the most spread-out measure with density matrix $\hat{\rho}$, see Section 7.1 for the definition). Among other things, we determine $\mathbb{P}_{\mathrm{gc}}$ in this paper (see Statements 3 and 4b in Section 3).
@@
(with Boltzmann’s constant $k_{\mathrm{B}}$) under the constraint that the expected energy is fixed, $\operatorname{tr}(\hat{\rho}\,\hat{H})=E$ (as well as, in the grand-canonical case, the expected particle numbers, $\operatorname{tr}(\hat{\rho}\,\hat{N}_{i})=n_{i}$). While this reasoning may suggest that $\hat{\rho}_{\mathrm{can}}$ ($\hat{\rho}_{\mathrm{gc}}$) represents the knowledge of an observer who knows not more about the state than $E$ (and $n_{i}$) (e.g., [24, 60], [1, Chap. 4], [2, Chap. 7]) the reasoning outlined above in terms of $S\cup B$ provides more [36]: a physical reason why the actual, physical state (of $S$) is given (approximately) by $\hat{\rho}_{\mathrm{can}}$ (or $\hat{\rho}_{\mathrm{gc}}$); see Section 4.1 for further discussion.
Also concerning the use of the micro-canonical distribution $u_{\mathrm{mc}}=u_{\mathbb{S}(\mathscr{H}_{\mathrm{mc}})}$ or density matrix $\hat{\rho}_{\mathrm{mc}}$, there are different attitudes (e.g., [2, Sec. 6.8]) in the literature: Sometimes its use is regarded as a “principle of equal a priori probabilities” (e.g., [51, 23, 35]) saying that unless we have further knowledge about $\Psi$ than that it lies in $\mathscr{H}_{\mathrm{mc}}$, it is a rational guess to assign each point on $\mathbb{S}(\mathscr{H}_{\mathrm{mc}})$ equal (subjective) probability. Our attitude is different: Since it is the nature of thermal equilibrium to behave like most $\Psi$, we find the thermal equilibrium behavior by studying the behavior of most $\Psi$, where “most” refers to $u_{\mathrm{mc}}$ (or, when appropriate, $u_{\mathrm{gmc}}$ over $\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$ as in (8)).
@@
$\hat{\rho}_{\mathrm{gG}}=Z^{-1}_{\mathrm{gG}}\;\exp\left(\sum_{k=1}^{K}\lambda_{k}\hat{Q}_{k}\right)\,$, (7)
which we also discuss as it plays, in fact, a key role. It represents the thermal equilibrium ensemble for which $\hat{Q}_{1},\ldots,\hat{Q}_{K}$ (and only those) are conserved macroscopic observables, [^1] The expression “macroscopic observable” is often intended to include the assumption that the eigenvalues of $\hat{Q}_{k}$ are separated by the resolution of macroscopic measurements; this is not intended here. assuming they commute with each other. (One of them usually is the Hamiltonian $\hat{H}$. Previous considerations in this direction can be found in [28, Sec. I.4].) Note that, as a consequence, there is a $K$-parameter family of thermal equilibrium states with parameters $\lambda_{1},\ldots,\lambda_{K}$ (which includes the canonical case with $K=1$, $\hat{Q}_{1}=\hat{H}$, and $\lambda_{1}=-\beta$). The generalized micro-canonical subspace $\mathscr{H}_{\mathrm{gmc}}$ for this situation is the one where the eigenvalues of each $\hat{Q}_{k}$ are restricted to an interval $Q_{k}-\Delta Q_{k},Q_{k}]$ that is short on the macroscopic scale but still has a high-dimensional spectral subspace $\mathds{1}_{[Q_{k}-\Delta Q_{k},Q_{k}]}(\hat{Q}_{k})\mathscr{H}$,
@@
We can give here a derivation in the case in which each of the $\hat{Q}_{k}$ is (approximately) extensive, that is, in which, for any spatial region $S$ and its complement $\overline{S}$,
$\hat{Q}_{k}\approx\hat{Q}_{k}^{S}\otimes\hat{I}^{\overline{S}}+\hat{I}^{S}\otimes\hat{Q}_{k}^{\overline{S}}\,.$ (13)
This is the first of our claims, see Statement 1a in Section 3. It includes the grand-canonical density matrix $\hat{\rho}_{\mathrm{gc}}$ as a special case (also for chemical equilibrium, see Section 2, but also for spatial equilibrium, where the conserved observables are simply $\hat{Q}_{k}=\hat{N}_{k}$).
@@ 2 Use of $\hat{\rho}_{\mathrm{gc}}$: Chemical Reactions
We answer both in this section. Our reasoning applies not only to chemical reactions but also to ionization and to elementary particle reactions.
@@
Remark 1.
There may be a (somewhat arbitrary) choice of the theoretical physicist involved in defining which states exactly count (say, in the example of (6)) as states of molecule $C$ as opposed to states of a molecule $A$ and a molecule $B$ [29, Sec. 3]. We assume in the following that such a choice has been made. In the case of ionization (say, $A$ = proton, $B$ = electron, $C$ = hydrogen atom), a natural choice [11] would be to count the bound states (from the discrete spectrum of the relative Hamiltonian) as atoms $C$ and the scattering states (from the continuous spectrum of the relative Hamiltonian) as $A+B$. $\diamond$
@@
we assume that the interaction term $\hat{V}$ is mostly negligibly small (namely, except during a chemical reaction) and contains non-zero transition elements corresponding to every reaction (14), but not for any transition that is not of the form (14). That is, we assume that the conserved quantities of (14) are the only macroscopic conserved observables. In detail, using (simple) tools from linear algebra, consider the $r$-dimensional real space $\mathcal{R}$ with axes labeled $n_{1},\ldots,n_{r}$, and let $\mathcal{L}$ be the subspace spanned by the $L$ vectors $(\nu_{\ell 1}-\tilde{\nu}_{\ell 1},\ldots,\nu_{\ell r}-\tilde{\nu}_{\ell r})$. Any change in particle numbers due to reactions of the form (14) lies in $\mathcal{L}$, any conserved linear combination of $n_{1},\ldots,n_{r}$ is orthogonal to $\mathcal{L}$. (In fact, the deeper reason we allow real rather than integer components in $\mathcal{R}$ is that we allow real coefficients in these linear combinations.) Choose a complete set $F_{1},\ldots,F_{K-1}$ of conserved linear combinations with
$K=\dim(\mathcal{L}^{\perp})+1=r+1-\dim\mathcal{L}$ (24)
(which is equal to r+1Lr+1-L if the $L$ vectors just mentioned are linearly independent). That is, choose a linear mapping (a $(K-1)\times r$-matrix) $F:\mathcal{R}\to\mathbb{R}^{K-1}$ with kernel
@@
$F$ has full rank, $\mathrm{rank}(F)=\dim\mathrm{image}(F)=K-1$, and for every $k=1,\ldots,K-1$, $F_{k}(n_{1},\ldots,n_{r})$ is conserved during every reaction (14). (Put differently, the rows $F_{1},\ldots,F_{K-1}$ of the matrix $F$ form a basis of $\mathcal{L}^{\perp}$.) As a consequence, the observables
@@
using (19); the function $(1/\beta)\log Z_{\mathrm{gc}}(\beta,\mu_{01},\ldots,\mu_{0r})$ is known as the grand potential (e.g., [1, (5.73)], [44, (2.7.13)]). Note that, while (37c) is an expression that one might have guessed right from the start, our derivation has achieved two things: First, provided reasons for why (37) is true, and second, provided a way of computing the $\mu_{0i}$ from the known quantities $n_{0i},\delta E_{\ell}$ by means of (35) and (10).
@@
In the example of (6), $E_{0A}$ and $E_{0B}$ can be chosen arbitrarily, $E_{0C}$ is determined by (20), (35) reduces to $\mu_{0C}=\mu_{0A}+\mu_{0B}+\delta E$, and (10) reduces to the three conditions (with $\hat{\rho}_{A}$ etc. as in (36))
@@
Remark 4.
The reaction rates depend on $\hat{V}$, but the equilibrium state and the $n_{\mathrm{eq},i}$ do not, as long as $\hat{V}$ is small. Related questions for non-small $\hat{V}$ have been studied in [7, 4]. $\diamond$
@@ 3 Main Claims
We now present our main statements. Apart from Statement 1b, the derivations of the statements are given below in the following sections. Let $\Lambda\subset\mathbb{R}^{3}$ denote the available volume of the quantum system considered and
$\overline{S}:=\Lambda\setminus S$ (39)
the spatial complement of a subset $S$ (not the closure of a set). We take $\hat{\rho}_{\mathrm{gmc}}$ as defined in (9) and $\hat{\rho}_{\mathrm{gG}}$ as defined in (7). For extensive $\hat{Q}_{k}$ as in (13), it follows that
$\hat{\rho}_{\mathrm{gG}}\approx\hat{\rho}_{\mathrm{gG}}^{S}\otimes\hat{\rho}_{\mathrm{gG}}^{\overline{S}}$ (40)
with $\hat{\rho}_{\mathrm{gG}}^{S}$ formed according to (7) from the $\hat{Q}_{k}^{S}$ and likewise for $\overline{S}$. This also entails that
$\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gG}}\approx\hat{\rho}_{\mathrm{gG}}^{S}\,.$ (41)
  • Statement 1a. (Generalized Gibbs Density Matrix) Let $\Lambda$ be a subset of $\mathbb{R}^{3}$, $\mathscr{H}=\mathscr{H}^{\Lambda}$ a Hilbert space, and for every $S\subseteq\Lambda$ let $\mathscr{H}^{S}$ be a Hilbert space such that for $S_{1}\subseteq S_{2}\subseteq\Lambda$, $\mathscr{H}^{S_{2}}$ factorizes into $\mathscr{H}^{S_{2}}=\mathscr{H}^{S_{1}}\otimes\mathscr{H}^{S_{2}\setminus S_{1}}$. Suppose the self-adjoint operators $\hat{Q}_{1},\ldots,\hat{Q}_{K}$ on $\mathscr{H}$ are bounded from below and commute with each other, and that each is (approximately) extensive as in (13). Then for $\lambda_{1},\ldots,\lambda_{K}$ satisfying (10), large $\dim\mathscr{H}_{\mathrm{gmc}}$, and small $S\subset\Lambda$,
    $\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}\approx\hat{\rho}_{\mathrm{gG}}^{S}\,.$ (42)
@@
    $\mathscr{H}_{1i}=\mathscr{H}_{1i}^{S}\oplus\mathscr{H}_{1i}^{\overline{S}}$ (44)
@@
    $\hat{H}_{1i}\approx\hat{H}_{1i}^{S}\oplus\hat{H}_{1i}^{\overline{S}}\,.$ (45)
    Suppose (14) for $\ell=1,\ldots,L$ are the (only) possible reactions (i.e., changes in particle numbers), and (21) holds. Choose any $F$ satisfying (25), and define $\hat{Q}_{k}$ by (27) (so that $\hat{H},\hat{Q}_{k}]=0$) and $\hat{Q}_{K}=\hat{H}$. Then for $\dim\mathscr{H}_{\mathrm{gmc}}\gg 1$ and $\mu_{0i}$ satisfying (10) and (35),
@@
    $\hat{\rho}_{\Psi}^{S}:=\operatorname{tr}^{\overline{S}}|\Psi\rangle\langle\Psi|\approx\hat{\rho}_{\mathrm{gG}}^{S}\,.$ (47)
@@
    $\hat{\rho}_{\Psi}^{S}=\operatorname{tr}^{\overline{S}}|\Psi\rangle\langle\Psi|\approx\hat{\rho}_{\mathrm{gc}}^{S}\,.$ (48)
@@
The conditional wave function can be thought of as what the collapsed wave function of $S$ would be after an ideal quantum measurement of the basis $b$ on $B$ [17, Footnote 2].
@@
We now answer the question, raised in the introduction, of what $\mathbb{P}_{\mathrm{gc}}$ is, the probability distribution of the wave function corresponding to the grand-canonical ensemble. We answer it by providing the typical distribution of the conditional wave function $\psi^{S}$. The answer is twofold, as it depends on how the basis $b$ is chosen. In general, $\psi^{S}$ (and its distribution) depends on the choice of $b$, and while $b$ can be chosen arbitrarily, one is often particularly interested in taking $b$ the position (or configuration) basis [5]. We provide two results assuming that $S$ is a small spatial region and $B=\overline{S}$: one result in which $b$ is a typical (i.e., random) ONB, and one in which $b$ is typical among those that diagonalize the conserved operators $\hat{Q}_{k}^{\overline{S}}$ (respectively, the particle number operators $\hat{N}_{i}^{\overline{S}}$). The first result (Statement 3) applies whenever $b$ is unrelated to the joint eigenbasis of $\hat{H}^{\overline{S}}$ and the $\hat{Q}_{k}^{\overline{S}}$ (respectively, $\hat{N}_{i}^{\overline{S}}$); the second (Statements 4a and 4b) whenever there is no simple relation between $b$ and $\hat{H}^{\overline{S}}$ other than that both $b$ and a suitable eigenbasis of $\hat{H}^{\overline{S}}$ diagonalize the $\hat{Q}_{k}^{\overline{S}}$ (the $\hat{N}_{i}^{\overline{S}}$)—this seems like a reasonable approximation for the position basis.
@@
  • Statement 3. ($\mathbb{P}_{\mathrm{gG}}$ and $\mathbb{P}_{\mathrm{gc}}$ for Typical ONB) In the setting of Statement 1a (or 1b) with $S$ a small spatial region, for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$ and most ONBs $b$ of $\mathscr{H}^{\overline{S}}$, the conditional wave function (49) has distribution
@@
  • Statement 4a. ($\mathbb{P}_{\mathrm{gG}}$ for ONB Diagonalizing $\hat{Q}_{k}^{\overline{S}}$) In the setting of Statement 1a, suppose that $S\subset\Lambda$ is a small region and that for every $k=1,\ldots,K-1$, $\Delta Q_{k}=0$ and $Q_{k}$ is an eigenvalue of $\hat{Q}_{k}$. Then for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$ and most ONBs $b$ of $\mathscr{H}^{\overline{S}}$ that diagonalize $\hat{Q}_{1}^{\overline{S}},\ldots,\hat{Q}_{K-1}^{\overline{S}}$ (but not $\hat{Q}^{\overline{S}}_{K}=\hat{H}^{\overline{S}}$),
@@
  • Statement 4b. ($\mathbb{P}_{\mathrm{gc}}$ for ONB Diagonalizing $\hat{N}_{i}^{\overline{S}}$) In the setting of Statement 1b, suppose that $S\subset\Lambda$ is a small region and that for every $k=1,\ldots,K-1$, $\Delta Q_{k}=0$ and $Q_{k}$ is an eigenvalue of $\hat{Q}_{k}$. Then for most $\Psi\in\mathbb{S}(\mathscr{H}_{\mathrm{gmc}})$ and most ONBs $b$ of $\mathscr{H}^{\overline{S}}$ that diagonalize $\hat{N}_{1}^{\overline{S}},\ldots,\hat{N}_{r}^{\overline{S}}$,
    $\mathscr{L}_{\psi^{S}}\approx\mathbb{P}_{\mathrm{gc}}:=\bigoplus_{n_{1}^{\overline{S}},...,n_{r}^{\overline{S}}}p_{n}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})~\mathrm{GAP}_{\hat{\rho}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})}\,$, (59)
@@
    $p_{n}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})=\operatorname{tr}(\hat{\rho}_{\mathrm{gmc}}\hat{P}_{n})~~\text{with}~~\hat{P}_{n}:=\hat{P}\Bigl(\hat{N}_{1}^{\overline{S}}=n_{1}^{\overline{S}},\ldots,\hat{N}_{r}^{\overline{S}}=n_{r}^{\overline{S}}\Bigr)$ (60)
    and
    $\hat{\rho}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})=\frac{1}{p_{n}(n_{1}^{\overline{S}},\ldots,n_{r}^{\overline{S}})}\operatorname{tr}^{\overline{S}}\bigl[\hat{P}_{n}\,\hat{\rho}_{\mathrm{gmc}}\,\hat{P}_{n}\bigr]\,.$ (61)
Remark 5.
If the general Gibbs principle is more generally valid than under the assumptions of Statement 1a (which is plausible as pointed out in Section 4.2), then the expression for $\mathbb{P}_{\mathrm{gc}}$ can be simplified further:
$\mathbb{P}_{\mathrm{gc}}\approx\bigoplus_{q_{1}^{S},...,q_{K-1}^{S}}p(q_{1}^{S},\ldots,q_{K-1}^{S})~\mathrm{GAP}_{\hat{\rho}^{S}(q_{1}^{S},\ldots,q_{K-1}^{S})}\,$, (62)
where
[…]
@@
$\diamond$
Statement 4b is actually not a special case of Statement 4a because only certain combinations of $\hat{N}_{i}$ are conserved as the $\hat{Q}_{k}$, and while in Statement 4a we considered bases diagonalizing the $\hat{Q}_{k}^{\overline{S}}$ (i.e., certain combinations of the $\hat{N}_{i}^{\overline{S}}$), we consider in Statement 4b bases diagonalizing all $\hat{N}_{i}^{\overline{S}}$. That is because we are interested in the position/configuration basis, which diagonalizes all $\hat{N}_{i}^{\overline{S}}$; on the other hand, Statement 4a is the simpler statement, and can be applicable to general $\hat{Q}_{k}$ that are not related to number operators.
@@
The next statement concerns the approach (time evolution) towards thermal equilibrium. As many earlier works such as [57, 37, 30, 15], we take “approach” to mean that for most $t\geq 0$ in the long run, the system is in thermal equilibrium. And we take “thermal equilibrium” to mean that $\hat{\rho}_{\Psi_{t}}^{S}\approx\hat{\rho}_{\mathrm{gG}}^{S}$ and $\mathscr{L}_{\psi^{S}}\approx\mathbb{P}_{\mathrm{gG}}$.